Lauren is planning to raise her batting average. When she set up her weekly practice schedule, her batting average was 250. Lauren intends to increase her batting average by 5 each week. Find a pattern and write a formula that will give her expected batting average after n weeks.
a1 =
d=
Using the formula, how many weeks will Lauren need to practice if she expects to earn a batting average of 300?

Answer by mananth(16946)

Lauren is planning to raise her batting average. When she set up her weekly practice schedule, her batting average was 250. Lauren intends to increase her batting average by 5 each week. Find a pattern and write a formula that will give her expected batting average after n weeks.
a1 =
d=
Using the formula, how many weeks will Lauren need to practice if she expects to earn a batting average of 300?
She starts her batting average with 250
She increases it by 5 every week
So we use the formula for arithmetic sequence
We can calculate average by the formula
tn= a+(n-1)d
Expected batting average = 350
a= 250, d=5 and n=? , tn= 350
350 = 250+(n-1)*5
350-250 = (n-1)*5
100/5 = n-1
20 =n-1
n=21
Lauren's expected batting average would be 350 after 21 weeks